# OAC: Output-adaptive Calibration for Accurate Post-training Quantization

### v1

@@ Abstract
Deployment of Large Language Models (LLMs) has major computational costs, due to their rapidly expanding size. Compression of LLMs reduces the memory footprint, latency, and energy required for their inference. Post-training Quantization (PTQ) techniques have been developed to compress LLMs while avoiding expensive re-training. Most PTQ approaches formulate the quantization error based on a layer-wise $\ell_{2}$ loss, ignoring the model output. Then, each layer is calibrated using its layer-wise Hessian to update the weights towards minimizing the $\ell_{2}$ quantization error. The Hessian is also used for detecting the most salient weights to quantization. Such PTQ approaches are prone to accuracy drop in low-precision quantization. We propose Output-adaptive Calibration (OAC) to incorporate the model output in the calibration process. We formulate the quantization error based on the distortion of the output cross-entropy loss. OAC approximates the output-adaptive Hessian for each layer under reasonable assumptions to reduce the computational complexity. The output-adaptive Hessians are used to update the weight matrices and detect the salient weights towards maintaining the model output. Our proposed method outperforms the state-of-the-art baselines such as SpQR and BiLLM, especially, at extreme low-precision (2-bit and binary) quantization.
@@ 1 Introduction
The development of Large Language Models (LLMs) has sparked a revolution in natural language processing, leading to remarkable advancements in various tasks, including but not limited to reasoning, question answering, text generation, and few-shot learning [30, 2, 42, 36, 37, 1]. LLMs comprise a huge number of parameters, exceeding tens and even hundreds of billions, which is considered one of the key factors to their success [2]. Their large size, however, is associated with major computational complexities and deployment barriers, especially, on resource-limited machines.
Post-training Quantization (PTQ) approaches have been introduced to overcome LLMs deployment challenges by reducing the precision of the model weights or activations while maintaining the performance without extra training [39, 40, 11, 23, 19, 8, 15]. PTQ techniques eliminate the necessity of Quantization-aware Training (QAT) methods [18, 24, 34, 9] that are computationally demanding for LLMs. PTQ methods are one of the main techniques for efficient deployment of LLMs, enabling near loss-less compression up to 4 bits [8]. However, extremely low-precision (2-bit or binary) PTQ is still an open challenge, which is the main focus of our proposed method.
To recover the performance of the quantized model, most novel PTQ methods [11, 3, 8, 15] perform a layer-wise calibration to reduce the quantization error. These methods measure the quantization error by the $\ell_{2}$ loss between the output of the original and quantized layers over a small calibration set as shown in Figure 1:a). We refer to such approaches as output-agnostic calibration because they only consider the output of the individual linear layers separately and ignore the model output. The response-agnostic methods compute the second derivative of the $\ell_{2}$ loss for the weights of each layer to formulate a layer-wise Hessian. The Hessian is used to measure the saliency of weights in order to detect the most salient ones as outliers [8, 19, 15]. Also, according to the Hessian, the weights are updated toward minimizing the $\ell_{2}$ quantization error [11, 3, 8, 15]. Given their output-agnostic nature, they are deemed inadequate to recover the performance of the original model at extreme low-precision quantization.
@@
OAC follows the layer-by-layer recipe to quantize and calibrate the linear layers of LLMs. Our proposed method, however, employs the second derivative of the output cross-entropy loss to compute the output-adaptive Hessian in contrast to the existing PTQ methods that use the Hessian of the layer-wise $\ell_{2}$ loss. We utilize the output-adaptive Hessian for updating the weights and measuring their saliency to accurately calibrate the model. The computation of the exact output-adaptive Hessian for large models is, however, computationally infeasible. To this end, we employ several techniques to reduce the computational complexity of approximating the Hessian.
Our proposed method, OAC, adopts the common assumptions in the PTQ literature [10, 11, 8], e.g. (i) the independence of linear layers, and (ii) the independence of the rows of the weight matrices. Assumption (i) is used for computation of the output-adaptive Hessian matrix of each layer, while assumption (ii) is used, in conjunction with the Fisher Information Identity, towards approximating the row-wise Hessians. Moreover, we formulate our quadratic optimization problem by aggregating the row-wise Hessian matrices to significantly reduce the memory footprint.
@@
  • We provide theoretical insights on how to approximate the Hessian and provide a thorough experimental evaluation of various tasks to support our proposed PTQ method.
@@ 2 Related Work
There are two main categories of quantization techniques notably known as Quantization-aware Training (QAT) and Post-training Quantization (PTQ). QAT methods perform the quantization and training simultaneously [25, 17, 21, 34]. Considering the computational costs of re-training large models, QAT techniques are quite costly. The viable alternative is employing PTQ techniques, which utilize accurate solvers to minimize the quantization error without further training. Many PTQ methods leverage calibration that slightly modifies the model weights to reduce the quantization error on a small calibration dataset [13].
@@
Among the existing PTQ approaches, AdaRound [28], OBQ [10], AdaQuant [16], and BRECQ [22] are designed and applied to small computer vision models with around 100 million parameters. However, applying these methods to LLMs with billions of parameters is computationally intensive.
@@
ZeroQuant [40], LLM.int8() [7], and SmoothQuant [39] are among the first techniques that investigate PTQ of LLMs. ZeroQuant explores the quantization granularity in addition to layer-wise knowledge distillation. LLM.int8() separates the outlier activations using a threshold while quantizing the model to INT8 format. SmoothQuant [39] reduces the activation quantization complexity by scaling the inputs of layers. Despite succeeding at 8-bit quantization of LLMs, ZeroQuant, SmoothQuant, and LLM.int8() perform poorly on extreme low-precision quantization such as 2-bit.
@@
Inspired by [14, 20], OPTQ [11] performs a column-wise calibration procedure on the weights, which enables more accurate 3- and 4-bit PTQ of LLMs using a reasonable budget on time and computational resources. AWQ [23] detects the salient weights and scales them to reduce the PTQ error. QuIP [3] uses a generalized version of the OPTQ’s update formula in addition to incoherence pre-processing of the weights and Hessians to reduce the accuracy drop at sub-4-bit PTQ. SpQR [8] improves the OPTQ’s calibration strategy with two major modifications to achieve near loss-less compression of LLMs up to 4-bits, (i) detecting and isolating the outliers by the Hessian (ii) performing a second round of quantization on the quantization parameters such as scales and zeros to reduce the average bit-width while keeping the outliers at FP32 format and using small group quantization.
@@
SqueezeLLM [19] investigates the non-uniform PTQ of LLMs using sensitivity-based K-means clustering without any calibration. To detect the salient weights, i.e. outliers, SqueezeLLM approximates the Hessian of each linear layer considering the output loss using the Fisher information identity. However, the non-uniform quantization is often associated with deployment challenges during the inference [13]. OmniQuant [34] introduced an efficient QAT method that freezes the model weights while learning the quantization parameters to achieve the accuracy of QAT methods with a light training recipe. PB-LLM [33] investigated partial binarization of LLMs through a two-stage quantization recipe, (i) performing an OPTQ-based PTQ method on the non-salient binarized weights, (ii) performing QAT on the quantized model to further recover the accuracy while the salient weights are frozen. BiLLM [15] developed a more accurate PTQ method for binarization of LLMs by identifying and structurally selecting salient weights. Furthermore, BiLLM minimizes the quantization error using a binary residual approximation strategy. BiLLM also employs a splitting search algorithm to group and quantize non-outlier weights based on their bell-shaped distribution.
@@
The following points are the advantages of our proposed method compared to the state-of-the-art PTQ methods. (i) Our method computes the output-adaptive Hessian to calibrate the weights considering the model output loss, while OPTQ, QuIP, SpQR, and BiLLM use the response-agnostic layer-wise Hessian. (ii) In contrast to SqueezeLLM and BRECQ, our method does not rely on the assumption that the output-adaptive Hessian is diagonal and achieves a more accurate approximation (see Section 4.4.2). Moreover, BRECQ cannot be scaled to LLMs due to its computational complexity. Non-uniform quantization of SqueezeLLM is associated with deployment challenges and fails to support extreme low-precision quantization.
@@ 3 Output-agnostic Calibration Background
In this section, we describe our notation as well as the response-agnostic calibration setting which is used by the status quo PTQ methods [10, 11, 3, 8, 15].
@@
Let $\mathbf{W}$ denote the weight matrix of the $l^{\mathrm{th}}$ linear layer. The quantization and calibration of $\mathbf{W}$ is modeled by adding a small update term, $\delta\mathbf{W}$, such that $\widehat{\mathbf{W}}=\mathbf{W}+\delta\mathbf{W}$. To accurately quantize the weights after training, it is necessary to define a loss function to measure the quantization error. In the response-agnostic setting [10, 11, 3, 8], the quantization error, $\varepsilon_{\ell_{2}}$, is measured by the $\ell_{2}$ loss between the output of the quantized and original layer
@@
where $\bar{\mathbf{H}}=\mathbb{E}_{\mathbf{x}^{(l)}}[\mathbf{x}^{(l)}\mathbf{x}^{(l)^{\top}}]$ is the response-agnostic Hessian and it is assumed that the rows of the weight matrix independently contribute to $\varepsilon_{\ell_{2}}$ [10].
@@
The most salient weights are labeled as outliers. However, various outlier mitigation techniques have been developed [8, 19, 15, 13]. Equation (3) is used to measure the saliency of each weight
@@ 4 Proposed Output-adaptive Calibration
where $\bar{\mathbf{H}}^{(\bm{\theta})}_{\mathrm{OAC}}=\mathbb{E}_{(\mathbf{x},\mathbf{y})}\Bigr{[}\frac{\partial^{2}\mathcal{L}_{\mathrm{CE}}}{\partial\bm{\theta}\partial\bm{\theta}^{\top}}\Bigr{]}\in\mathbb{R}^{D\times D}$ is the output-adaptive Hessian of the entire weights and $\bar{\mathbf{g}}=\mathbb{E}_{(\mathbf{x},\mathbf{y})}\Bigr{[}\frac{\partial\mathcal{L}_{\mathrm{CE}}}{\partial\bm{\theta}}\Bigr{]}\simeq 0$ since the model is trained [28]. For LLMs, the computation of $\bar{\mathbf{H}}^{(\bm{\theta})}_{\mathrm{OAC}}$ is infeasible due to its huge size $\mathcal{O}(D^{2})$. In Sections 4.1, 4.2, and 4.3, we illustrate how our proposed approach circumvents the computation of exact $\bar{\mathbf{H}}^{(\bm{\theta})}_{\mathrm{OAC}}$.
@@ 4.1 Cross-layer Independence
Considering the dimensions of linear layers in LLMs, $\bar{\mathbf{H}}_{\mathrm{OAC}}$ is a huge matrix $\mathcal{O}(d_{\mathrm{row}}^{2}d_{\mathrm{col}}^{2})$ with memory-intensive computation. Some existing methods, such as [19, 22], ignore all of the inter-element correlations and assume that all of the weights are independent. Such methods approximate the Hessian of each layer by a diagonal matrix comprising the diagonal elements of the Fisher information matrix. However, this over-restrictive assumption leads to accuracy drop [14]. To achieve a more accurate Hessian approximation, we relax this assumption as described in Section 4.2.
@@ 5 OAC Pipeline
Figure 3: A demonstration of the OAC steps for 2-bit PTQ of LLMs. 1) The transformer blocks are iteratively selected for calibration. 2) The final outputs for the calibration samples are generated. 3) The generated outputs are compared with the ground truth outputs to compute the loss and gradients. 4) The output-adaptive Hessians of linear layers inside the block are approximated. 5) For each linear layer, the outliers are detected and isolated using $\widehat{\mathbf{H}}_{\mathrm{OAC}}$. 6) Column-wise calibration is performed to reduce the quantization error. 7) The quantization scales and zeros go through a second round of quantization to reduce the average bit width. Steps 5, 6, and 7 are integrated into our method from [8].
Our proposed method, OAC, comprises two main phases for PTQ of LLMs weights. (i) Computing the output-adaptive Hessian of each weight matrix. (ii) Calibrating each weight matrix using its output-adaptive Hessian. To develop a complete PTQ pipeline, we integrate a Hessian-based calibration technique with our proposed method.
@@
OAC uses equation (11) to approximate the Hessian of each weight matrix. Therefore, it requires computing the gradients of each weight matrix for all calibration samples. We propose computing the gradients of the linear layers inside each transformer block simultaneously, while other transformer blocks are frozen. This avoids repeating the costly output generation and backpropagation for all linear layers. As shown in Figure 3:a), the transformer blocks are iteratively selected. For each block, the model outputs are generated using the calibration samples. Following the calculation of the cross-entropy loss, OAC employs backpropagation to compute the gradients of the weights associated with the linear layers within the block. Then, the $\widehat{\mathbf{H}}_{\mathrm{OAC}}$ of the linear layers within the block are separately approximated based on equation (11) and are used in the next phase of our proposed PTQ method.
@@
To develop OAC for accurate 2-bit PTQ of LLMs, the following steps from SpQR [8] are integrated into our method. The salient weights are detected and isolated using equation (3) while $\bar{\mathbf{H}}$ is replaced by our proposed output-adaptive Hessian $\widehat{\mathbf{H}}_{\mathrm{OAC}}$. Then, the column-wise updating process is performed to reduce the distortion of the cross-entropy loss. Likewise, the update term is computed by replacing $\widehat{\mathbf{H}}_{\mathrm{OAC}}$ in equation (2). Finally, the scales and zeros are quantized to reduce the average bit hence enabling smaller group quantization and keeping more outliers in the FP32 format. Figure 3 illustrates the complete steps of OAC, implemented for 2-bit PTQ of LLMs. Furthermore, to show the effectiveness of our proposed output-adaptive Hessian computation in binary PTQ, we integrated the $\widehat{\mathbf{H}}_{\mathrm{OAC}}$ to the calibration procedure of BiLLM [15, Algorithm 1]. experimental results presented in Section 6 show that our proposed approach surpasses both BiLLM and SpQR in the PTQ of LLMs.

### v2

@@ Abstract
Deployment of Large Language Models (LLMs) has major computational costs, due to their rapidly expanding size. Compression of LLMs reduces the memory footprint, latency, and energy required for their inference. Post-training Quantization (PTQ) techniques have been developed to compress LLMs while avoiding expensive re-training. Most PTQ approaches formulate the quantization error based on a layer-wise Euclidean loss, ignoring the model output. Then, each layer is calibrated using its layer-wise Hessian to update the weights towards minimizing the quantization error. The Hessian is also used for detecting the most salient weights to quantization. Such PTQ approaches are prone to accuracy drop in low-precision quantization. We propose Output-adaptive Calibration (OAC) to incorporate the model output in the calibration process. We formulate the quantization error based on the distortion of the output cross-entropy loss. OAC approximates the output-adaptive Hessian for each layer under reasonable assumptions to reduce the computational complexity. The output-adaptive Hessians are used to update the weight matrices and detect the salient weights towards maintaining the model output. Our proposed method outperforms the state-of-the-art baselines such as SpQR and BiLLM, especially, at extreme low-precision (2-bit and binary) quantization.
@@ 1 Introduction
The development of Large Language Models (LLMs) has sparked a revolution in natural language processing, leading to remarkable advancements in various tasks, including but not limited to reasoning, question answering, text generation, and few-shot learning (Radford et al., 2019; Brown et al., 2020; Zhang et al., 2022; Touvron et al., 2023a, b; Achiam et al., 2023). LLMs comprise a huge number of parameters, exceeding tens and even hundreds of billions, which is considered one of the key factors to their success (Brown et al., 2020). Their large size, however, is associated with major computational complexities and deployment barriers, especially, on resource-limited machines.
Post-training Quantization (PTQ) approaches have been introduced to overcome LLMs deployment challenges by reducing the precision of the model weights or activations while maintaining the performance without extra training (Xiao et al., 2023; Yao et al., 2022; Frantar et al., 2023; Lin et al., 2024; Kim et al., 2024; Dettmers et al., 2024; Huang et al., 2024). PTQ techniques eliminate the necessity of Quantization-aware Training (QAT) methods (Kim et al., 2023; Liu et al., 2024; Shao et al., 2024; Du et al., 2024) that are computationally demanding for LLMs. PTQ methods are one of the main techniques for efficient deployment of LLMs, enabling near loss-less compression up to 4 bits (Dettmers et al., 2024). However, extremely low-precision (2-bit or binary) PTQ is still an open challenge, which is the main focus of our proposed method.
To recover the performance of the quantized model, most novel PTQ methods (Frantar et al., 2023; Chee et al., 2023; Dettmers et al., 2024; Huang et al., 2024) perform a layer-wise calibration to reduce the quantization error. These methods measure the quantization error by the $\ell_{2}$ loss between the output of the original and quantized layers over a small calibration set as shown in Figure 1:a). We refer to such approaches as output-agnostic calibration because they only consider the output of the individual linear layers and ignore the model output. The output-agnostic methods compute the second derivative of the $\ell_{2}$ loss for the weights of each layer to formulate a layer-wise Hessian. The Hessian is used to measure the saliency of weights in order to detect the most salient ones as outliers (Dettmers et al., 2024; Kim et al., 2024; Huang et al., 2024). Also, according to the Hessian, the weights are updated toward minimizing the $\ell_{2}$ error (Frantar et al., 2023; Chee et al., 2023; Dettmers et al., 2024; Huang et al., 2024). Given their output-agnostic nature, they are deemed inadequate to recover the performance of the original model at extreme low-precision quantization.
@@
OAC follows the layer-by-layer recipe to quantize and calibrate the linear layers of LLMs. Our proposed method, however, employs the second derivative of the output cross-entropy loss to compute the output-adaptive Hessian in contrast to the existing PTQ methods that use the Hessian of the layer-wise $\ell_{2}$ loss. We utilize the output-adaptive Hessian for updating the weights and measuring their saliency to accurately calibrate the model. The computation of the exact output-adaptive Hessian for large models is, however, computationally infeasible. To this end, we employ several techniques to reduce the computational complexity of approximating the output-adaptive Hessian.
Our proposed method, OAC, adopts the common assumptions in the PTQ literature (Frantar and Alistarh, 2022; Frantar et al., 2023; Dettmers et al., 2024), e.g. (i) the independence of linear layers, and (ii) the independence of the rows of the weight matrices. Assumption (i) is used for computation of the output-adaptive Hessian matrix of each layer, while assumption (ii) is used, in conjunction with the Fisher Information Identity, towards approximating the row-wise Hessians. Moreover, we formulate our quadratic optimization problem by aggregating the row-wise Hessian matrices to significantly reduce the memory footprint.
@@
  • We provide theoretical insights on how to approximate the Hessian and provide a thorough experimental evaluation of various tasks to support our proposed method.
@@ 2 Related Work
There are two main categories of quantization techniques notably known as Quantization-aware Training (QAT) and Post-training Quantization (PTQ). QAT methods perform the quantization and training simultaneously (Liu et al., 2024; Jacob et al., 2018; Li et al., 2017; Shao et al., 2024). Considering the computational costs of re-training large models, QAT techniques are quite costly. The viable alternative is employing PTQ techniques, which utilize accurate solvers to minimize the quantization error without further training. Many PTQ methods leverage calibration that slightly modifies the model weights to reduce the quantization error on a small calibration set (Gholami et al., 2022).
@@
Among the existing PTQ methods, AdaRound (Nagel et al., 2020), OBQ (Frantar and Alistarh, 2022), AdaQuant (Hubara et al., 2021), and BRECQ (Li et al., 2021) are designed and applied to small computer vision models with around 100 million parameters. However, applying these methods to LLMs with billions of parameters is computationally intensive.
@@
ZeroQuant (Yao et al., 2022), LLM.int8() (Dettmers et al., 2022), and SmoothQuant (Xiao et al., 2023) are among the first techniques that investigate PTQ of LLMs. ZeroQuant explores the quantization granularity in addition to layer-wise knowledge distillation. LLM.int8() separates the outlier activations using a threshold while quantizing the model to INT8 format. SmoothQuant (Xiao et al., 2023) reduces the activation quantization complexity by scaling the inputs of layers. Despite succeeding at 8-bit quantization of LLMs, ZeroQuant, SmoothQuant, and LLM.int8() perform poorly on extreme low-precision quantization such as 2-bit.
@@
Inspired by (Hassibi and Stork, 1992; LeCun et al., 1989), OPTQ (Frantar et al., 2023) performs a column-wise calibration procedure on the weights, which enables more accurate 3- and 4-bit PTQ of LLMs using a reasonable budget on time and computational resources. AWQ (Lin et al., 2024) detects the salient weights and scales them to reduce the PTQ error. QuIP (Chee et al., 2023) uses a generalized version of the OPTQ update formula in addition to incoherence pre-processing of the weights and Hessians to reduce the accuracy drop at sub-4-bit PTQ. QuIP# (Tseng et al., 2024) enhances QuIP by randomized Hadamard transform (Halko et al., 2011), vector quantization, and fine-tuning. SpQR (Dettmers et al., 2024) improves the OPTQ calibration strategy with two major modifications to achieve near loss-less compression of LLMs up to 4-bits, (i) detecting and isolating the outliers by the Hessian (ii) performing a second round of quantization on the quantization parameters such as scales and zeros to reduce the average bit-width while keeping the outliers at FP32 format and using small group quantization.
@@
SqueezeLLM (Kim et al., 2024) investigates the non-uniform PTQ of LLMs using sensitivity-based K-means clustering without any calibration. To detect the salient weights, SqueezeLLM approximates the Hessian of each linear layer considering the output loss using the Fisher Information Identity. However, the non-uniform quantization is often associated with deployment challenges during the inference (Gholami et al., 2022). OmniQuant (Shao et al., 2024) introduced an efficient QAT method that freezes the model weights while learning the quantization parameters to achieve the accuracy of QAT methods with a light training recipe. Orthogonal to our work, AQLM (Egiazarian et al., 2024) uses Additive Quantization (Babenko and Lempitsky, 2014) as well as fine-tuning to overcome low-precision quantization. PB-LLM (Yuan et al., 2024) investigated partial binarization of LLMs through a two-stage quantization recipe, (i) performing an OPTQ-based PTQ method on the non-salient binarized weights, (ii) performing QAT on the quantized model to further recover the accuracy while the salient weights are frozen. BiLLM (Huang et al., 2024) developed a more accurate PTQ method for the binarization of LLMs by identifying and structurally selecting salient weights. Furthermore, BiLLM minimizes the quantization error using a binary residual approximation strategy. BiLLM also employs splitting search to group and quantizes non-outlier weights based on their bell-shaped distribution.
Our proposed method has the following advantages compared to the state-of-the-art PTQ methods.
  • Our method computes the output-adaptive Hessian to calibrate the weights considering the model output cross-entropy loss, while OPTQ, QuIP, QuIP#, SpQR, and BiLLM use the response-agnostic layer-wise Hessian.
@@
  • In contrast to SqueezeLLM and BRECQ, our method does not rely on the assumption that the output-adaptive Hessian is diagonal, and achieves a more accurate approximation. Moreover, BRECQ cannot be scaled to LLMs due to its computational complexity. Non-uniform quantization of SqueezeLLM is associated with deployment challenges and also fails to support extreme low-precision quantization.
@@ 3 Output-agnostic Calibration Background
In this section, we describe our notation as well as the response-agnostic calibration setting which is used by the status quo PTQ methods (Frantar and Alistarh, 2022; Frantar et al., 2023; Chee et al., 2023; Dettmers et al., 2024; Huang et al., 2024).
@@
Let $\mathbf{W}$ denote the weight matrix of the $l^{\mathrm{th}}$ linear layer. The quantization and calibration of $\mathbf{W}$ is modeled by adding a small update term, $\delta\mathbf{W}$, such that $\widehat{\mathbf{W}}=\mathbf{W}+\delta\mathbf{W}$. To accurately quantize the weights after training, it is necessary to define a loss function to measure the quantization error. In the response-agnostic setting (Frantar and Alistarh, 2022; Frantar et al., 2023; Chee et al., 2023; Dettmers et al., 2024), the quantization error, $\varepsilon_{\ell_{2}}$, is measured by the $\ell_{2}$ loss between the output of the quantized and original layer
@@
where $\bar{\mathbf{H}}=\mathbb{E}_{\mathbf{x}^{(l)}}[\mathbf{x}^{(l)}\mathbf{x}^{(l)^{\top}}]$ is the response-agnostic Hessian and it is assumed that the rows of the weight matrix independently contribute to $\varepsilon_{\ell_{2}}$ (Frantar and Alistarh, 2022).
@@
The most salient weights are labeled as outliers. However, various outlier mitigation techniques have been developed (Dettmers et al., 2024; Kim et al., 2024; Huang et al., 2024; Gholami et al., 2022). Equation (3) is used to measure the saliency of each weight
@@ 4 Proposed Output-adaptive Calibration
where $\bar{\mathbf{H}}^{(\bm{\theta})}_{\mathrm{OAC}}=\mathbb{E}_{(\mathbf{x},\mathbf{y})}\Bigr{[}\frac{\partial^{2}\mathcal{L}_{\mathrm{CE}}}{\partial\bm{\theta}\partial\bm{\theta}^{\top}}\Bigr{]}\in\mathbb{R}^{D\times D}$ is the output-adaptive Hessian of the entire weights and $\bar{\mathbf{g}}=\mathbb{E}_{(\mathbf{x},\mathbf{y})}\Bigr{[}\frac{\partial\mathcal{L}_{\mathrm{CE}}}{\partial\bm{\theta}}\Bigr{]}\simeq 0$ since the model is trained (Nagel et al., 2020). For LLMs, the computation of $\bar{\mathbf{H}}^{(\bm{\theta})}_{\mathrm{OAC}}$ is infeasible due to its huge size $\mathcal{O}(D^{2})$. In Sections 4.1, 4.2, and 4.3, we illustrate how our proposed approach circumvents the computation of exact $\bar{\mathbf{H}}^{(\bm{\theta})}_{\mathrm{OAC}}$.
@@ 4.1 Cross-layer Independence
Considering the dimensions of linear layers in LLMs, $\bar{\mathbf{H}}_{\mathrm{OAC}}$ is a huge matrix $\mathcal{O}(d_{\mathrm{row}}^{2}d_{\mathrm{col}}^{2})$ with memory-intensive computation. Some existing methods, such as (Kim et al., 2024; Li et al., 2021), ignore all of the inter-element correlations and assume that all of the weights are independent. Such methods approximate the Hessian of each layer by a diagonal matrix comprising the diagonal elements of the Fisher information matrix. However, this over-restrictive assumption leads to accuracy drop (Hassibi and Stork, 1992). To achieve a more accurate Hessian approximation, we relax this assumption as described in Section 4.2.
@@ 5 OAC Pipeline
Algorithm 1 OAC Pipeline
0:  model, $N$ data samples
1:  for $j=1$ to model.num_layers do
[…]
3:     for $i=1$ to $N$ do
4:        loss = Cross-Entropy(model(data[$i$]), data[$i$])
5:        loss.backward()
6:        for layer $\in$ block.linear_layers do
[…]
8:           $\mathbf{G}[i]=\nicefrac{{\partial\text{loss}}}{{\partial\mathbf{W}}}$
9:           $\widehat{\mathbf{H}}_{\text{OAC}}+=\frac{1}{N}\mathbf{G}^{\top}[i]\mathbf{G}[i]$
10:        end for
11:     end for## Phase 2: Calibration by $\widehat{\mathbf{H}}_{\mathrm{OAC}}$#
12:     for layer $\in$ block.linear_layers do
[…]
14:     end for
15:  end for
Our proposed method, OAC, comprises two main phases for PTQ of LLMs weights. (i) Computing the output-adaptive Hessian of each weight matrix. (ii) Calibrating each weight matrix using its output-adaptive Hessian. To develop a complete PTQ pipeline, we integrate a Hessian-based calibration technique with our proposed method. The OAC pipeline is described in Algorithm 1.
@@
OAC uses equation (11) to approximate the Hessian of each weight matrix. Therefore, it requires computing the gradients of each weight matrix for all calibration samples. We propose computing the gradients of the linear layers inside each transformer block simultaneously, while other transformer blocks are frozen. This avoids repeating the costly output generation and backpropagation for all linear layers. As described in 1, the transformer blocks are iteratively selected. For each block, the model outputs are generated using the calibration samples. Following the calculation of the cross-entropy loss, OAC employs backpropagation to compute the gradients of the weights associated with the linear layers within the block. Then, the $\widehat{\mathbf{H}}_{\mathrm{OAC}}$ of the linear layers within the block are separately approximated based on equation (11) and are used in the next phase of our proposed PTQ method.
Figure 3: A demonstration of the OAC steps for 2-bit PTQ of LLMs. 1) The transformer blocks are iteratively selected for calibration. 2) The final outputs for the calibration samples are generated. 3) The generated outputs are compared with the ground truth outputs to compute the loss and gradients. 4) The output-adaptive Hessians of linear layers inside the block are approximated. 5) For each linear layer, the outliers are detected and isolated using $\widehat{\mathbf{H}}_{\mathrm{OAC}}$. 6) Column-wise calibration is performed to reduce the quantization error. 7) The quantization scales and zeros go through a second round of quantization to reduce the average bit width. Steps 5, 6, and 7 are integrated into our method from (Dettmers et al., 2024).
@@
To develop OAC for accurate 2-bit PTQ of LLMs, the following steps from SpQR (Dettmers et al., 2024) are integrated into our method. The salient weights are detected and isolated using equation (3) while $\bar{\mathbf{H}}$ is replaced by our proposed output-adaptive Hessian $\widehat{\mathbf{H}}_{\mathrm{OAC}}$. Then, the column-wise updating process is performed to reduce the distortion of the cross-entropy loss. Likewise, the update term is computed by replacing $\widehat{\mathbf{H}}_{\mathrm{OAC}}$ in equation (2). Finally, the scales and zeros are quantized to reduce the average bit hence enabling smaller group quantization and keeping more outliers in the FP32 format. Figure 3 illustrates the complete steps of OAC, implemented for 2-bit PTQ of LLMs. Furthermore, to show the effectiveness of our proposed output-adaptive Hessian computation in binary PTQ, we integrated the $\widehat{\mathbf{H}}_{\mathrm{OAC}}$ to the calibration procedure of BiLLM (Huang et al., 2024, Algorithm 1). Experimental results presented in Section 6 show that our proposed approach surpasses both BiLLM and SpQR in the PTQ of LLMs.